A triangle with an area of 120 cm2 has a height of 10 cm. What is the triangle's base measurement?
Answer by nerdybill(7384)

A triangle with an area of 120 cm2 has a height of 10 cm. What is the triangle's base measurement?
.
Since we know area of a triangle is defined as:
area = (1/2)bh
where
b is the length of the base
h is the height
.
From the information provided by the problem, we get:
120 = (1/2)b(10)
120 = b(5)
120/5 = b
24 cm = b (base)